# A modulated nonlinear system: breathers and time crystals

### v1

@@ II Physical system and dynamical equations
III Effect of space-time modulation on the linear dispersion
@@
For a finite system with $N$ units and periodic boundary conditions the possible values of $q$ are $q_{m}=2\pi m/N$, with $m=0,\dots,N-1$, this includes also negative values of the wavevector $q=-q^{\prime}$, with $q^{\prime}>0$, as any wave number $q$ is equivalent to $q\pm 2\pi n$. Also, any set of numbers $m^{\prime}=m+n$, where $n$ is any integer lead to an equivalent basis, a convenient one, centered around $q=0$ is $N/2,\dots,N/2-1$ if $N$ is even or $(N-1)/2,\dots,(N-1)/2$ if $N$ is odd. An alternative basis in real space is given by $\cos(qn-\omega t)$ and $\sin(qn-\omega t)$, but the exponential basis functions are more convenient analytically.
We first describe the effect of introducing the modulation term with $\delta>0$ small as observed in numerical simulations. Then we will obtain approximately the new dispersion relations, for several cases and the general case.
Using initial uniformly random conditions for $u_{n}$ and letting the system evolve to thermal equilibrium and performing the 2D fast Fourier transform (2DFFT) we obtain the energy density plot and the $\omega-q$ spectrum shown in Fig. 2 :
Figure 2: Numerical simulations after thermalization with modulation frequency and wave number $\Omega=0.02$ and $h=0.5$ : (Top:) Energy density plot (Bottom) 2D FFT.
@@
III.A Dispersion bands modification with only space modulation
@@
$u_{n,t}$ $=\exp(\textrm{i}q_{l}n)\exp(\textrm{i}q_{r}n)\exp(-\textrm{i}\omega_{l,r}t)$
@@
Figure 3: (Top:) Numerical dispersion bands for $h=2\pi/3$; (Bottom:) Theoretical degeneracy raising of the dispersion bands due to space modulation. Note that the Brillouin zone is $0,2\pi/3]$. See the text for explanation.
III.B Effect of only time modulation on the phonon bands
When $h=0$, the effect of time modulation is to split the phonon band in several parallel bands separated by a frequency of $\Omega$. The number of bands diminishes with increased frequency and eventually the bands disappear or are not visible. Extra bands appear for frequency below $\omega_{0}=1$ although they are very weak when approaching $\Omega=\omega_{0}$. Frequencies multiples of $\omega_{0}$ lead to a divergence of the system. Frequencies inside the phonon band excites the modes with that frequency.
In the following we will use the notation and theory exposed in Appendix A, taking into account the changes brought about by the time dependence of the on-site potential.
III.C Breakup of the time-invariance: first approximation
@@
It is invariant under translations in time $t\rightarrow t+\Delta t$, where $\Delta t$ is the sampling interval, which could be arbitrary, in principle. Let us choose $\Delta t=T_{m}/M$, with $M$ an integer, that is $\Omega=2\pi/T_{m}=M2\pi/\Delta t$. The addition of the term $\delta\cos(\Omega t)u_{n}$ to the rhs of (21) breaks this invariance. The modulated system is invariant only under larger translation of time $T_{m}=M\Delta t$, multiples of $\Delta t$. There is a breaking of the symmetry to a smaller symmetry.
@@
$+\frac{\displaystyle\delta}{\displaystyle 2}(A_{m+1}+A_{m-1})\,$, ($2$)
@@
$(\omega+m\Omega)^{2}A_{m}=\bar{\omega}_{q}^{2}A_{m}+\frac{\displaystyle\delta}{\displaystyle 2}(A_{m+1}+A_{m-1})\,$, (23)
@@
Therefore, the main effect of the symmetry breaking is the apparition of bands separated by the frequency $\Omega$ almost parallel to the unperturbed dispersion band (UDB). However, the intensity of those bands tend to be smaller the further apart from the UDB. It is necessary to plot contours of small intensity in order to observe them.
@@
Figure 4: (Top) Contour plot of the 2DFFT for $\delta=0.05$, $\Omega=0.1$, $\kappa=0.1$ (Bottom) Identical parameters except $\Omega=0.01$
III.D Dispersion bands for space-time modulation
@@
Assuming that $\delta$ is small with respect to $\Omega$, we can obtain the independent solutions where all the $A_{k}=0$, except one $A_{m}$ each time, so it correspond to a single harmonic with $\omega$:
@@
Figure 5: Contour plot of the FFT2 for $\delta=0.05$, $\Omega=0.05$, $\kappa=0.1$ and the theoretical dispersion relation lines: Note that there is an area to the right where the harmonics are very close and become mixed.
IV Breathers with space-time modulation
In this section we study the conditions for breather existence in a modulate system, first reviewing the theory of exact breathers in Archilla et al. (2019) and adapting it to space-time modulation and to just time modulation.
IV.A Theory of exact traveling breathers
In this section we follow closely the theory developed in Ref. Archilla et al. (2019). Let us suppose that we have a space time modulated system with phase $\phi=hn-\Omega t$ and velocity $V_{m}=h/\Omega$. An exact breather will be a solution of the form
$u(n,t)=f(n-V_{b}t,\omega_{i}t)\,$, (29)
where $\omega_{F}=2\pi/T_{F}$ is the fundamental frequency, $T_{F}$ is the fundamental period, defined as the minimal time for which $u$ reproduces itself displaced a distance $s=V_{b}T_{F}$ called the step. The function $f$ is a localized function of its first argument and a $2\pi$ periodic function of the second. The frequency $\omega_{i}$ is the moving frame or internal frequency for short.
$\mathrm{with}\quad\omega_{L}=qV_{b}+m\omega_{F}\,.$ (30)
IV.B Application to space-time harmonically modulated systems
Therefore:
$h(n+s)-\Omega(t+T_{F})=hn-\Omega t+2\pi m_{m}\,\rightarrow$
The immediate consequence for only time modulation is:
Figure 6: (Top) For $\Omega=2\omega_{\text{b}}$, profile of the stable breather. (Bottom) Coordinates and energy evolution in a breather period $T_{b}$.
Figure 7: For $\Omega=2\omega_{\text{b}}$ Floquet multipliers (left) for the non-autonomous system and (right) for the extended autonomous system. See text.
IV.C Breather obtention in the time-modulated system
The most direct way is working on the real space with the Newton and shooting method. Trying some different localized initial conditions and observing an approximate periodical behavior with period close to $T_{b}$, the intended one, the initial conditions are used as a seed for the Newton method where initial conditions corresponding to the exact $T_{b}$ periodic behavior. Exact means with a precision of $10^{-15}$, that is, very close, to the computer machine precision. The Newton method, a numerical application of the implicit function theorem, can be used as the system has no time invariance due to the time-periodic term in the potential, and a solution is unique in its neighborhood. The precision can be refined more by working in the space of the frequencies, if we consider time-reversible solutions, they can be written as $u_{n}(t)=\sum_{k=-k_{m}}^{k_{m}}z_{k,n}\exp(k\omega_{\text{b}}t)$. Due to $u_{n}(t)$ being real and time-symmetric, $z_{k,n}^{*}=z_{-k,n}$ and $z_{k,n}=z_{-k,n}$, and real. The $u_{n}(t)=z_{0,n}+\sum_{k=1}^{k_{m}}2z_{k,n}\cos(k\omega_{\text{b}}t)$, being 2zk,n2z{k,n} the inverse cosine fast Fourier cosine transform (idfc). Then, the Newton method can be coded in the space of the $k_{m}+1$ FFT components, finding the fftc components $z_{n,k}$. The value of $m=15$ provides excellent results.
IV.D Stability of breathers in the time-modulated system
Let as suppose that the breather solution for a time crystals with $T_{b}=n_{T}T_{m}$, with $n_{T}$ an integer larger than 1., the $T_{m}$-periodic system is also $T_{b}$ periodic, and to analyze the stability we can construct the Floquet matrix integrating a small perturbation of the initial coordinates and momentum and observing the perturbation after $T=T_{b}$. For each changed coordinate we obtain a column of the Floquet matrix $F$. Its eigenvalues lie at the unit circle if the system is stable. For autonomous systems symplectic systems Aubry (2006) there are always two eigenvalues of $F$ at +1+1, corresponding to $u_{t}=\dot{u}$, the phase mode, as it represents that if $u(t)$ is a $T_{b}$ periodic solution, $u(t+dt)$ is also a solution, and the growth mode, indicating that around a solution there is another one with slightly different amplitude.
Those two eigenvalues do not generally appear for the non-autonomous system as can be seen in Fig. 7-left for $\Omega=2\omega_{\text{b}}$. The explanation is that if we write the dynamical equations (9) in a condensed form as:
$N_{n}\xi$ $=\ddot{\xi}_{n}-L_{n}\xi$
$\ddot{\dot{u}}_{b,n}-L_{n}\dot{u}_{b,n}$
where the last term does not appear in (33).
However, it can be recovered within an autonomous extended system that includes the autonomous one as it is shown below. For the different factors $n_{T}$, in $\Omega=n_{T}\omega_{\text{b}}$, we obtain different results that we will present later.
Figure 8: (Top) Floquet eigenvalues for a breather with $\Omega=3\omega_{\text{b}}$ and $\omega_{\text{b}}=0.98$ (Bottom) Fourier spectra of the same breather embedded in a noisy system with $N=160$ cantilevers.
IV.E Extended autonomous symplectic system
The extended system reveals a hidden symplectic structure of the autonomous system with the implication that stability corresponds to all the Floquet multipliers being in the unit circle, but without the double multiplier of +1, which is recovered with the extended system. Figure 7-right shows these multipliers for $\Omega=2\omega_{\text{b}}$.
V Time crystals: breathers with time modulation and $\Omega=n_{T}\omega_{\text{b}}$
We consider the nonlinear dynamical system (9) with two main cases, $\omega_{\text{b}}=\Omega/n_{T}$ ($T_{\text{b}}=n_{T}T_{m}$), and $\omega_{\text{b}}=n_{T}\Omega$ ($T_{\text{b}}=T_{m}/n_{T}$), $n_{T}$ an integer. The first one correspond to time crystals Yao and Nayak (2018); Wilczek (2019), while the latter has no specific name. Although, according to the deduction in Sec. IV, fractional frequencies are also possible, we will not consider them in this article. A condition for proper time crystals is the no production of entropy, which happens here as the energy is repeated after $T_{b}$ with no energy dissipation as shown in Fig. 7.
In this section we consider time crystals and leave the condition of $\Omega<\omega_{\text{b}}$ for the next section.
For $\Omega=2\omega_{\text{b}}$, and changing slightly $\omega_{\text{b}}$, there is an interval where the two phonon bands coincide and then the system becomes stimulated by $\Omega$ and the variables diverge. A dissipation term can be added, but as this a very special case, we prefer not to modify the system and simply avoid those combination of frequencies.
V.A Period doubling $\Omega=2\omega_{\text{b}}$
In the figures we can see the profile of breather and neighbors, also the evolution of the energy, which is $T_{m}$-periodic and therefore also $T_{b}$-periodic. There is no net gain or loss of energy every period, which indicates that the solution is a true time-crystal. The Floquet eigenvalues are also pictured. The eigenvalues corresponding to the phonons form a pair of complex conjugate bands width frequencies $\omega_{i}=\omega_{\text{b}}+\theta_{i}/T_{b}$, as they have the same Krein signature Aubry (2006) they form a structurally stable subset. There is also a close but separate eigenvalue with a localized eigenvector, and thus corresponding to the breather solution. It has the opposite Krein signature of the phonons, meaning it can collide with the phonon with closest frequency leaving the unit circle, which indicates that the breather is unstable and another solution exists, most likely and extended one, corresponding to the merger of the breather with the upper phonon mode.
Figure 9: (Top) Profile and Floquet eigenvalues for a breather with $\Omega=\omega_{\text{b}}/10$ and $\omega_{\text{b}}=0.98$ (Bottom) Evolution of the coordinates and the AC current.
V.B Cases $\omega_{\text{b}}=3\Omega$ to $\omega_{\text{b}}=10\Omega$
For $\omega_{\text{b}}=3\Omega$ the breather is slightly unstable with two real multipliers very close to +1. In spite of that simulations can be done to simulate the breather evolution with relatively long time. The Floquet multipliers are represented in Fig. 8 and the Fourier spectrum in a system with noise and $N=160$ oscillators are shown.
For $\omega_{\text{b}}=n_{T}\Omega$, width $n_{T}=1\dots 10$, we have constructed breathers, for $n_{T}=3,4,8$ breathers are unstable, while the rest are stable, therefore there is no apparent pattern. For example, for $n_{T}=3$, there are two two real eigenvalues slightly separated from +1+1, but, in spite of that, the breather has a long life.
VI Breathers with $\Omega=\omega_{\text{b}}/n_{T}$
A peculiarity of this case is that the stability of the breather and the Floquet matrix have to be calculated for a time equal to the modulation period $T_{m}=n_{T}T_{\text{b}}$ as the coordinates, and velocities have to repeat but also the forces which depend on time an only repeat after $T_{m}$. Therefore, the phonon band appear with Floquet exponents multiplied by $n_{T}$ and spread out they are given $\theta_{q}=\omega_{q}T_{m}=n_{T}\omega_{q}T_{b}$ for a phonon of frequency $\omega_{q}$.
In this section we separate the study of the Sievers-Takeno mode Sievers and Takeno (1988b) or single breather, from the Page mode Page (1990), or double breather. Values $n_{T}=1,2,\dots,32$ are considered and a breather obtained. For $n_{T}=32,\dots,40$ phonobreathers appear as explained in Sec. VI.B.
Figure 10: Case $\Omega=\omega_{\text{b}}/20$ and $\omega_{\text{b}}=0.98$ (Top) Frequency-momenta representation with scaled frequency. To the left for $N=16$ sites, and to the right for $N=160$ and noise added. See text. (Bottom) Spectra of the Fourier intensities in arbitrary units with respect to the scaled frequency. See text.
VI.A Sievers-Takeno mode: single ILM with $\omega_{\text{b}}=n_{T}\Omega$
In out system, the unmodulated system dispersion relation is given by $\omega^{2}=\omega_{0}^{2}+2\kappa(1-\cos(q))$, with a minimum frequency $\omega_{0}=1$. As the nonlinear part of the on-site potential is soft, because the dominant term is $o(3)$, stationary breathers or ILMs will have a frequency $\omega_{\text{b}}$ below the minimum frequency $\omega_{0}$. As demonstrated in Sect. IV, for a time-modulated system, the modulating frequency $\Omega$, is commnesurate with $\omega_{\text{b}}$, and in this subsection, we consider the case $\Omega=\omega_{\text{b}}/n_{T}$. By using the shooting method, we tried and obtained exact ILMs for many values of $n_{T}$, as 1, 2, …40. We present here only some results with $\omega_{\text{b}}=0.98$, that is, very close to the minimum frequency $\omega_{0}=1$ for the unmodulated system.
We construct the ILM in a small lattice with $N=16$ cantilevers, and check its stability by constructing the monodromy matrix and obtaining the Floquet multipliers and exponents Aubry (2006). Then we embed the initial positions and within a larger lattice with $N=160$ particles with some small random noise velocities, let it evolve and perform the 2D FFT to check the results. An example $\Omega=\omega_{\text{b}}/10$ with $\omega_{\text{b}}=0.98$ can be seen in Fig. 9. The top panel shows the profile at $t=0$ and the Floquet eigenvalues, showing its stability. The bottom panel illustrates the evolution of the variables for a period of the AC current.
We choose the case $\Omega=\omega_{\text{b}}/20$ with $\omega_{\text{b}}=0.98$, to illustrate the frequency-momenta representation as it is more easily seen. The top panel of Fig. 10-top shows the theoretical spectra $\omega=\pm m\Omega+\sqrt{\omega_{0}^{2}+2\kappa(1-\cos(q)}$. The horizontal dotted line is the scaled physical frequency $\omega_{0}/\omega_{0,\text{phys}}$ of an isolated oscillator without modulation for reference (See Sect. II). To the right it is shown the numerical spectra obtained for the larger lattice with some noise added that makes the dispersion relation to appear. The figure is obtained by adding up the intensities for all the particles at the same time.
We can see that also that frequencies $\omega_{\text{b}}\pm\Omega$ appear displaced $\omega_{0}-\omega_{\text{b}}=1-0.98=0.02$ above and below the bottom of the secondary dispersion curves. Horizontal lines for the ILM are due to the extreme localization in $u_{n}$ and consequent dispersion in $q$.
Figure  10-bottom shows the spectra intensity with respect to the frequency, where the phonon bands and the extra ILM bands can also be observed. The upper and lower secondary phonon and ILM bands are symmetric, there are more secondary bands for $m>1$ with diminishing intensities.
Figure 11: (Top) For $\Omega<\omega_{\text{b}}/33$, a phonobreather appears as a background to the ILM. (Bottom) Fourier spectra. See text.
VI.B Single ILM with phonobreather background for $\omega_{\text{b}}=n_{T}\Omega$
Although the obtention of ILM for increasing $n_{T}$ is possible, as $\Omega$ becomes smaller, the upper secondary ILM band $\omega_{\text{b}}+\Omega$ collides with the upper frequency of the phonon band $\omega_{\text{max}}=\sqrt{\omega_{0}+2\kappa(1-\cos(\pi))}=\sqrt{\omega_{0}+4\kappa}=\omega_{\text{b}}+\Omega$, or $\Omega=\omega_{\text{max}}-\omega_{\text{b}}$ and $\omega_{\text{b}}/\Omega\simeq 32.3$. Then for $n_{t}>30$ the upper phonons are excited and a phonobreather Archilla et al. (2003) with $q=\pi$ profile appears as the background of the ILM, with smaller but comparable amplitude. A phonobreather is a periodic solution on a nonlinear system with homogeneous amplitude. It is not a phonon as it is not a solution of the linearized system and has a significant amplitude. It can also be described as a nonlinear extended wave.
The profile of the breather with phonobreather background can be seen in Fig. 11-(Top) and at the bottom of the same figure, the Fourier spectra, where the colliding of the top of the phonon bands can be seen merging with the ILM bands of upper order, while in Fig. 10-bottom they are well separated.
Figure 12: (Top) For $\Omega<\omega_{\text{b}}/10$, profile of the unstable in-phase Page mode. (Bottom) The stable anti-phase Page mode with the same parameters and modulating frequency.
VI.C Page modes with $\omega_{\text{b}}=n_{T}\Omega$
A Page modePage (1990); Flach and Gorbach (2005) is a double breather or ILM with two sites vibrating with the same amplitude, it can be bond-symmetric if both sites oscillate in-phase (iP), or bond-antisymmetric if they oscillate in anti-phase (aP). The iP mode is unstable and the aP is stable for an unmodulated Klein Gordon system with soft potential Archilla et al. (2003). Both two modes appear also in the modulated system with the same stability properties. Both mode profiles and Floquet multipliers are represented in Fig. 12 for $\omega_{\text{b}}=10\Omega$. The 2DFFT of the in-phase mode shows the appearance of another breather band closer to the bottom of the phonon band, which correspond to a the unstable mode. Similar phenomena to the single ILM occur, i.e., formation of replicas of the phonon band and ILM band displaced $\pm\Omega$ in the frequency space and the excitation of a phonobreather background when the secondary ILM band interacts with another replica of the phonon band just below.
VII Conclusions
We have analyzed the peculiarities of the stability analysis for time-modulated systems with the construction of and extended symplectic system. Time-crystal breathers, that is, breathers with a period multiple of the modulating system are observed and their properties analyzed.
For the modulating frequency smaller than the breather’s one, we have found single breathers and double breathers, both symmetric and anti-symmetric. When the modulating frequency becomes small a phonobreather background appear.
The study reveals that this relatively simple system provides a great variety of breather solutions which may also appears in other space-time modulated materials also known as dynamic metamaterialsWu et al. (2021).
Acknowledgements
Funding
YD acknowledges support from JSPS Kakenhi (C) No. 24K14978.
Appendix A Notation
The following relations are useful and easy to obtain:
$=\sum_{k=0}^{N-1}\frac{1}{\sqrt{N}}\exp(-\textrm{i}\frac{\displaystyle 2\pi nk}{\displaystyle N})F_{k}(u)\,$, (39)
A.A Functions of position and time
Suppose we have a function $u=$ of the particle number $n$ and time $t$, its value at a particle and time $u_{n}(t)$, are the components of the function $u$ represented as the ket $|u\rangle$ in the basis with elements $|n,t\rangle$, i.e. $\langle n,t|u\rangle=u_{n}(t)$. In principle, $t$ is a continuous variable but, in practice, it is a discrete one obtaining by the sampling of numerical integration. So, if there are $N_{t}$ time samples, $t_{l}=l\Delta t$, with $l=0:N_{t}-1$ and $\Delta t$ the sampling interval. We change the basis notation to $|n,l\rangle$, with $\langle n,t|u\rangle=u_{n}(t_{l})=u_{n}(l\Delta t)$ and
$|u\rangle=\sum_{n=1}^{N}\sum_{l=0}^{N_{t}-1}u_{n}(t_{l})|n,l\rangle\,.$
The basis $|k,s\rangle$ is also orthonormal as
Therefore,
and then:

### v2

@@ A.1 Functions of position and time
Appendix B Deduction of the effect of space-time modulation on the linear dispersion
@@
For a finite system with $N$ units and periodic boundary conditions the possible values of $q$ are $q_{m}=2\pi m/N$, with $m=0,\dots,N-1$, this includes also negative values of the wavevector $q=-q^{\prime}$, with $q^{\prime}>0$, as any wave number $q$ is equivalent to $q\pm 2\pi n$. Also, any set of numbers $m^{\prime}=m+n$, where $n$ is any integer lead to an equivalent basis, a convenient one, centered around $q=0$ is $N/2,\dots,N/2-1$ if $N$ is even or $(N-1)/2,\dots,(N-1)/2$ if $N$ is odd. An alternative basis in real space is given[$\in[-0.5,0.5]$by $\cos(qn-\omega t)$ and $\sin(qn-\omega t)$, but the exponential basis functions a,rt=5000t=5000, for example, when the system is supposed to be near or at ore convenient analytically.
We first describe the effect of introducing the modulation term with $\delta>0$ small as observed in numerical simulations. Then we will obtain approximately the new dispersion relations for several cases.
B.1 Obtention of the numerical dispersion curves
We obtain numerically the dispersion bands with the following procedure: First we consider a relatively large system, with, for example, $N=128$ oscillators, and use as initial conditions a random distribution of positions for the variables $u_{n}$ in [-0.5,0.5], while the velocities are kept at zero. Equally, the random perturbation can be added to the momenta or both the momenta and coordinates without changing the outcome. Second, we leave the linear system (7) to evolve a relatively long time, $t=5000$, for example, when the system is supposed to be near or at thermal equilibrium and therefore all the linear modes or phonons are in principle excited. Then, we use the final positions and velocities, to run a simulation of similar length. Then, we use the results to perform the two dimensional fast fourier transform XTFFT on the variable $u_{n}(t)$, that is,
$F(q_{k},\omega_{r})=\frac{1}{\sqrt{N_{t}N}}\sum_{n=0}^{N-1}\sum_{l=0}^{N_{t}-1}\exp(\textrm{i}[q_{k}n-\omega_{r}t_{l}])u_{n}(t_{l})\,$,
with $t_{l}=\frac{\displaystyle l}{\displaystyle N_{t}}t_{f}$, $\omega_{r}=\frac{\displaystyle r}{\displaystyle N_{t}}\frac{\displaystyle 2\pi}{\displaystyle\Delta t}$, for $l,r=0\dots N_{t}-1$ and $q_{k}=\frac{\displaystyle k}{\displaystyle N}2\pi$, for $k=0\dots N-1$. The sampling interval $\Delta t=\frac{\displaystyle t_{f}}{\displaystyle N_{t}}$ is larger that the integration step to avoid very large matrices as the maximum frequency $2\pi/\Delta t$ does not need to be very large because the frequencies are of the order or $\omega_{0}$ and $\Omega$. On the other hand the frequency precision $\Delta\omega=\frac{\displaystyle 2\pi}{\displaystyle N_{t}\Delta t}=\frac{\displaystyle 2\pi}{\displaystyle t_{f}}$, is inversely proportional to $t_{f}$, therefore a large value of it is convenient, both to obtain precision for the frequency and for the results to be generic enough.
The intensities are defined as $I(\omega_{i},q_{n})=|F(\omega_{i},q_{n})|^{2}$, normalized to unity. To visualize the dispersion curves, we represent a number of lines in contour plots, typically ten, to get a clear picture. However, all the modes do not have the same probability to occur, the density of states may be small, and theoretically existing modes have small intensities and are difficult to discriminate from the noise background, those dispersion lines are weak. Presently, we do not have a means to calculate theoretically those intensities.
This method is changed to check that the results do not depends on the details, the random distribution can be applied only to momenta or to both coordinates and momenta; we can change the thermalization time, simulation time, sampling time interval, number of particles, etc. This method will also be used with a superposed breather in other sections.
The result of the simulation after thermalization are also used to obtain energy density plots. An example of both is shown in Fig. 2.
@@ B.1 Obtention of the numerical dispersion curves
B.2 Dispersion bands modification with only space modulation
@@
$u_{n,t}$ $=\exp(\textrm{i}q_{l}n)\exp(\textrm{i}q_{r}n)\exp(-\textrm{i}\omega_{l,r}t)$ (30)
@@
Figure 15: (Top:) Numerical dispersion bands for $h=2\pi/3$; (Bottom:) Theoretical degeneracy raising of the dispersion bands due to space modulation. Note that the Brillouin zone is $0,2\pi/3]$. See the text for explanation. Parameters: $\kappa=0.10.055794$, $\delta=0.05$, $\Omega=0$ .
B.3 Effect of only time modulation on the phonon bands
When $h=0$, the effect of time modulation is to split the phonon band in several parallel bands separated by a frequency of $\Omega$. They are obtained numerically as explained in App. B.1. Some bands are weak, meaning that their intensity in the XTFFT is small, presumably because their density of states is small. Some becomes so weak that they can not be seen into the background noise. The number of bands diminishes with increased frequency and eventually the bands disappear or are not visible. Extra bands appear for frequency below $\omega_{0}=1$ although they are very weak when approaching $\Omega=\omega_{0}$. Frequencies multiples of $\omega_{0}$ lead to a divergence of the system. Frequencies inside the phonon band excites the modes with that frequency.
These bands can be obtained analytically using the breakup of the time-invariance brought about by the time modulation. For that, we will use the notation and theory treated in Appendix A, taking into account the changes brought about by the time dependence of the on-site potential.
@@ B.3 Effect of only time modulation on the phonon bands
It is invariant under translations in time $t\rightarrow t+\Delta t$, where $\Delta t$ is the sampling interval, which could be arbitrary, in principle. Let us choose $\Delta t=T_{m}/M$, with $M$ an integer, that is $\Omega=2\pi/T_{m}=M2\pi/\Delta t$. The addition of the term $\delta\cos(\Omega t)u_{n}$ to the rhs of (42) breaks this invariance. The modulated system is invariant only under larger translation of time $T_{m}=M\Delta t$, multiples of $\Delta t$. There is a breaking of the symmetry: the system is no longer invariant under the time shifts $t\rightarrow t+\Delta t$ but only under a subgroup of time shifts $t\rightarrow t+T_{m}=t+M\Delta t$.
@@
$+\frac{\displaystyle\delta}{\displaystyle 2}(A_{m+1}+A_{m-1})\,$, (43)
@@
$(\omega+m\Omega)^{2}A_{m}=\bar{\omega}_{q}^{2}A_{m}+\frac{\displaystyle\delta}{\displaystyle 2}(A_{m+1}+A_{m-1})\,$, ($4$)
@@
Therefore, the main effect of the symmetry breaking is the appearance of bands separated by the frequency $\Omega$ almost parallel to the unperturbed dispersion band (UDB). However, the intensity of those bands tend to be smaller the further apart from the UDB. It is necessary to plot contours of small intensity in order to observe them.
@@
B.4 Dispersion bands for space-time modulation
@@
Assuming that $\delta$ is small with respect to $\Omega$, we can obtain the independent solutions where all the $A_{k}=0$, except one $A_{m}$ each time, so it corresponds to a single harmonic with $\omega$: